24. The inflation rate in the U.S. is 3%, while the inflation rate in Japan is 10%. The current exchange rate for the Japanese yen () is $0.0075. After supply and demand for the Japanese yen has adjusted in the manner suggested by purchasing power parity, the new exchange rate for the yen will be:
Answer #1

Answer - The new exchange rate for yen will be $0.0070

Reason -

New exchange rate = ( 1 + US rate ) ÷ (1 + Japan rate) * current exchange rate

= (1 + 0.03) ÷ ( 1 + 0.10) * 0.0075

= 0.9363636 * 0.0075

= 0.007022727 or $0.0070 ( rounded off)